# Predicting Census Survey Response Rates with Parsimonious Additive Models and Structured Interactions

### v4

@@ 2 Statistical Models and Methodology
We now discuss the statistical models we pursue in this work. Section 2.1 gives an overview of AMs with nonlinear main effects and pairwise interactions, along with optimization formulations associated with the estimation procedures. Section 2.2 presents our new models to incorporate sparsity in the main and interaction effects.
@@ 2.1 Smooth additive models with pairwise interactions
Given data $\{(y_{i},\mathbf{x}_{i})\}_{1}^{n}$, our key objective is to learn a multivariate conditional mean function $f(\mathbf{x}):=\mathbb{E}(y|\mathbf{x})$, where $f:\mathbb{R}^{p}\mapsto\mathbb{R}$ is an unknown smooth function (Wahba, 1990). It is well known that such functions become difficult to estimate even for moderately high $p$ due to the curse of dimensionality – therefore we will focus on a smaller class of models corresponding to additive structures (Hastie and Tibshirani, 1987; Stone, 1986). A popular approach considered in the literature estimates a nonparametric additive model containing main-effects only, with $f(\mathbf{x})=\sum_{j=1}^{p}f_{j}(x_{j})$, where each $f_{j}$ is an unknown univariate smooth function of the $j$-th coordinate in $\mathbf{x}$, namely $x_{j}$. In various applications, however, nonparametric AMs based on main effects alone may not lead to a sufficiently rich representation for predicting the outcome of interest: including interaction terms can lead to better predictive models while remaining interpretable to a practitioner (Hastie and Tibshirani, 1987). AMs with pairwise interactions are a useful tool in applied statistical modeling with various applications in medical sciences and healthcare, e-commerce applications, recommender system problems, and sentiment analysis, among others (Hastie, Tibshirani and Wainwright, 2015).
The need for interactions. Exploratory analysis on US Census Planning Database Tract dataset suggests that interaction effects across features can lead to improved predictions of the survey self-response rate. For example, there is interaction between the percentages of “People who do not speak English well” and “Renters” in the area. In areas with a relatively high concentration of poor English speakers (e.g, $\geq 5.4\%$), the self-response rate decreases on an average by $0.33\%$ for a unit increase in the percentage of renters. On the other hand, when the concentration of poor English speakers is relatively low (e.g, $\leq 1.6\%$), the self-response rate decreases at the rate of $0.21\%$. Similarly, there is a strong interaction effect between covariates “Single-unit households” and “Household moved in 2010 or later” in terms of predicting the low self-response rate. Indeed, Erdman and Bates (2016) note the importance of incorporating interaction effects when predicting the self-response rate, although the paper does not pursue statistical modeling involving interactions.
@@ 2.2 Parsimonious models via $\ell_{0}$-penalization
Here we study $\ell_{0}$-type estimators that limit the number of components in the additive models introduced earlier.
@@ 2.2.1 Sparse pairwise interactions
While a significant body of work has been devoted to studying sparsity in the context of linear models, sparsity in nonlinear models has received relatively less attention. Interestingly, we observe that the notion of parsimony is linked to the model being used and changes, for example, depending on whether we use a linear interaction model of the form $\mathbb{E}(y|\mathbf{x})=\sum_{j}x_{j}\beta_{j}+\sum_{j<k}x_{j}x_{k}\beta_{jk}$ or model (1) that has nonlinear components. For motivation, we consider Figure 3, which presents our findings on a 2019 US Census Bureau Planning Database dataset with $40$ covariates [^7] These covariates were selected based on discussions with researchers at the US Census Bureau.. These features include covariates used for the low-response-score (Erdman and Bates, 2014), important covariates highlighted in Appendix C of the 2019 US Census Bureau Planning Database Documentation (US Census Bureau, 2019), plus some additional covariates capturing internet penetration and urbanization. Specifically, we observe that:
@@
  • The above nonparametric additive model leads to better test predictions compared to its linear model counterpart.
  • The best (based on validation tuning with prediction error) model with linear main and interaction effects contains $37$ main and $555$ interaction effects. On the other hand, using nonparametric AMs, we need a total of 36 nonlinear main and interaction effects to obtain a similar test prediction performance. The nonlinear model also uses fewer covariates compared to the linear model counterpart.
The above observations suggest that the nonlinear model with sparse interactions has advantages over its linear model counterpart for predicting survey self-response rates. As discussed in Section 5, the above findings generally carry over to the expanded dataset with a larger number of features.
@@
Figure 3: Sparsity pattern of the main and interaction effects presented in a $p\times p$ matrix: a black square on the diagonal indicates the presence of a main effect, and an off-diagonal black square indicates the presence of an interaction effect in the joint model. [Left] Panel illustrates the sparsity pattern of a Lasso model with main and interaction effects. There are 37 main and 555 interaction effects in the optimal model. [Right] Panel illustrates the sparsity pattern of a nonlinear AM with main and interaction effects, i.e., model (8). There are 8 main and 92 interaction effects in the optimal model. Model (8) has prediction performance similar to the Lasso model, with only 3 main and 33 interaction effects. Nonlinear models lead to much more compact and hence, easier to interpret models compared to linear models. Both models were trained on a 2019 US Census Bureau Planning Database dataset (predicting the tract-level self-response rate) with $p=40$ covariates and $74,000$ observations.
@@
to simplify the expressions, and propose an estimator that optimizes the following problem:
@@
Related Work. In the special case of (8), when no interactions are present, several convex relaxations of the group $\ell_{0}$-penalty have been studied (Ravikumar et al., 2009; Zhao and Liu, 2012; Meier, van de Geer and Buhlmann, 2009; Buhlmann and van de Geer, 2011) [^10] In terms of existing implementations, R package SAM presents specialized algorithms for a convex relaxation of (8) without interactions. SAM however would not run on the dataset we consider here.. Hazimeh, Mazumder and Radchenko (2023) consider $\ell_{0}$-formulations for the setting without interaction effects and demonstrate the merits of using $\ell_{0}$-based formulations over convex relaxation-based approaches. Despite the appeal of estimator (8), computational challenges appear to be a key limiting factor in exploring this model in practice. In Section 2.4, we present a new algorithm for problem (8).
@@ 2.3 Statistical Theory
In this section, we explore statistical error bounds of our proposed estimator in the deterministic design setting. Technical details, assumptions, and proofs of the established results are provided in the supplementary material.
@@
Theorem 2.1 presented below derives a general non-asymptotic oracle prediction error bound, comparing the performance of our estimator to that of sparse approximations to the true regression function. Bounds of this type have been established for sparse nonparametric AMs with main effects; however, we are not aware of similar existing bounds for nonparametric models with pairwise interactions. The existing work has focused mainly on the Lasso-based estimators, which encourage sparsity in the main-effects by penalizing the magnitudes of the functional components (see Meier, van de Geer and Buhlmann, 2009; Tan and Zhang, 2019, and the references therein). The use of $\ell_{1}$-based relaxations, in lieu of $\ell_{0}$-penalization, to induce sparsity in the main-effects can lead to unwanted shrinkage, which may interfere with variable selection. The approach in Hazimeh, Mazumder and Radchenko (2023) controls the number of main-effects directly, demonstrating the benefits of $\ell_{0}$-regularization both theoretically and empirically. Our analysis accounts for the interaction effects and focuses on the $\ell_{0}$-penalized formulation, which poses additional theoretical challenges relative to the $\ell_{0}$-constrained formulation. To our knowledge, the error bounds established here for sparse nonparametric AMs with pairwise interactions are novel.
@@
[…]
@@
We note that a popular approach in the literature, even in the linear setting, has been to assume a sparse underlying model for the data. In contrast, Theorem 2.1 does not assume a ground truth sparse additive model, allowing for model misspecification and incorporating the approximation error into the bound. This feature is especially appealing for our application of interest where the ground truth is not available. The following corollary complements the general result in Theorem 2.1 by establishing a non-asymptotic prediction error bound for the proposed approach in the case where the additive model is correctly specified.
@@ 2.4 Efficient Computations at Scale
We present specialized algorithms for obtaining solutions to Problem (8). Our approach scales to the problem-size of the Census application, with $n\approx 10^{5}$ and $p\approx 500$, which poses formidable computational challenges due to the presence of approximately $10^{5}$ interaction effects [^11] Note that, using 25 knots for every component, this leads to estimating around 2.5 million basis coefficients.. To obtain good solutions at scale, we use techniques inspired by first order methods in continuous optimization (Nesterov, 2003) and a careful exploitation of the problem-structure. A high-level summary is presented below, with the details relegated to the Supplement.
@@
State-of-the-art approaches. To appreciate the computational challenges of sparse nonlinear AMs with interactions, and nonparametric AMs in general, we provide a few examples of the problem instances that can be handled by current state-of-the-art algorithms with publicly available implementations. Current implementations based on R package SAM (Zhao and Liu, 2012), the stepwise GAM function in R package step.gam (which performs greedy variable selection), and Python package pyGAM, take on the order of days to run and/or face numerical difficulties for instances with $n\approx 10^{5}$ to obtain a single solution without interactions. Furthermore, pyGAM is unable to do automated variable selection. Wood et al. (2017) present an interesting approach for AMs that scales to large $n$ settings (see R package mgcv) but doesn’t appear to perform automated variable selection in the presence of a large number of features – Wood et al. (2017) report instances containing fewer than $20$ pre-specified main and interaction effects. EBM (Nori et al., 2019) is a state-of-the-art computational approach for AMs, but does not allow for feature selection as all the main effects are included in the estimated model. Additionally, the variant of EBM that selects interaction effects, results in many interaction effects, leading to sub-optimal support recovery – see Sections 4 and 5 for an illustration. The GAMI-Net method (Yang, Zhang and Sudjianto, 2021) is computationally expensive. Additionally, since it uses a multi-layered NN for every interaction effect it requires 100×100$\times$100 × more parameters when compared with the approaches we consider.
@@
Algorithms for sparse nonlinear interactions: Problem (8). We represent the main and interaction effects as linear combinations of cubic spline basis functions (see the Supplement for the specific details). In particular, we let $\mathbf{f}_{j}=\mathbf{B}_{j}\boldsymbol{\beta}_{j}$, where $\mathbf{B}_{j}\in\mathbb{R}^{n\times K_{j}}$ is the model matrix and $\boldsymbol{\beta}_{j}\in\mathbb{R}^{K_{j}}$ is the vector of coefficients for each main effect component. Similarly, we let $\mathbf{f}_{j,k}=\mathbf{B}_{j,k}\boldsymbol{\theta}_{j,k}$, where $\mathbf{B}_{j,k}\in\mathbb{R}^{n\times K_{j,k}}$ is the model matrix and $\boldsymbol{\theta}_{j,k}\in\mathbb{R}^{K_{j,k}}$ is the vector of coefficients for each interaction effect component. Writing $\boldsymbol{\beta}$ for the vector obtained by stacking together the coefficients $\boldsymbol{\beta}_{j}$, $j\in[p]$ for the main-effects, and defining the vector $\boldsymbol{\theta}$ for the interaction effects analogously, we express the objective function in (4) as follows:
@@
Here, $\boldsymbol{S}_{j}=\boldsymbol{D}_{j}^{T}\boldsymbol{D}_{j}$ and $\boldsymbol{S}_{j,k}=(\boldsymbol{D}_{j}^{T}\boldsymbol{D}_{j})\otimes\boldsymbol{I}_{k}+\boldsymbol{I}_{j}\otimes(\boldsymbol{D}_{k}^{T}\boldsymbol{D}_{k})$ are the smoothness penalty matrices for the main effects and the interaction components, respectively (further details are provided in the Supplement). For convenience, we use the same smoothness penalty $\lambda_{1}$ for both the main and the interaction effects, though in general they may be taken to be different. The above representation leads to the following form of the optimization problem (8):
@@
We note that the indicator functions in the above equation are applied to vectors of basis coefficients corresponding to particular main or interaction effects. Hence, for example, when $\mathbb{1}\bigl{[}\boldsymbol{\beta}_{j}\neq\boldsymbol{0}\bigr{]}$ equals zero, the entire main effect $\mathbf{f}_{j}$ is also zero. As mentioned earlier, (12) can be expressed as a MIP and solved for small-to-moderate scale problems using modern MIP solvers. However, given the problem-sizes of interest and the fact that we seek to compute solutions to (12) for a family of tuning parameters, we discuss faster alternatives. We note that the objective in Problem (12) is a sum of a smooth convex loss function and a discontinuous regularizer separable across the components $\{\boldsymbol{\beta}_{j}\}$ and $\{\boldsymbol{\theta}_{i,k}\}$. Motivated by the strong empirical performance of cyclical coordinate descent (CD) methods (Wright, 2015) in $\ell_{0}$-penalized linear regression (Hazimeh and Mazumder, 2020a), we explore block CD methods to obtain fast approximate solutions for the nonparametric setting with interactions (12). For convergence guarantees of this procedure, see Hazimeh and Mazumder (2020a) and references therein.
@@
The block minimization problems (13) and (14) can be solved in closed-form as discussed in Supplement. These block CD updates need to be paired with several computational devices in the form of active set updates, cached matrix factorizations, and warm-starts, among others. We draw inspiration from similar strategies used in CD-based procedures for sparse linear regression (Hazimeh and Mazumder, 2020a; Friedman, Hastie and Tibshirani, 2010), and extend them to our problem. See Supplement Section S1.3 for more details.
CD methods have old roots in optimization dating back to the foundations of the discipline: see for example, the review paper Wright (2015). CD methods have close links with the well known Gauss-Seidel method (for solving linear/nonlinear systems of equations)—similar methods have been extensively used in additive models (Hastie and Tibshirani, 1990), where they are referred to as backfitting. CD schemes (including their block variants) arising from the optimization literature can be efficiently applied to optimization problems involving sparsity constraints: such problems arise in fitting additive models with sparsity-inducing penalties (Ravikumar et al., 2009; Hastie, Tibshirani and Wainwright, 2015; Hazimeh, Mazumder and Radchenko, 2023). We note that there are other successful additive model fitting approaches, for example, smooth backfitting (Mammen, Linton and Nielsen, 1999; Mammen and Park, 2006), which have excellent theoretical properties. These works focus on the additive model setup without sparsity—extending these approaches to our setting of large-scale sparse additive models in a computationally efficient fashion is an interesting direction for future research.

### v5

@@ 2 Statistical Models and Methodology
We now discuss the statistical models we pursue in this work. Section 2.1 gives an overview of AMs with nonlinear main effects and pairwise interactions, along with optimization formulations associated with the estimation procedures. Section 2.2 presents our proposed models to incorporate sparsity in the main and interaction effects.
@@ 2.1 Smooth additive models with pairwise interactions
Given data $\{(y_{i},\mathbf{x}_{i})\}_{1}^{n}$, our key objective is to learn a multivariate conditional mean function $f(\mathbf{x}):=\mathbb{E}(y|\mathbf{x})$, where $f:\mathbb{R}^{p}\mapsto\mathbb{R}$ is an unknown smooth function (Wahba, 1990). It is well known that such functions become difficult to estimate even for moderately high $p$ due to the curse of dimensionality – therefore, we will focus on a smaller class of models corresponding to additive structures (Hastie and Tibshirani, 1987; Stone, 1986). A popular approach considered in the literature estimates a nonparametric additive model containing main effects only, with $f(\mathbf{x})=\sum_{j=1}^{p}f_{j}(x_{j})$, where each $f_{j}$ is an unknown univariate smooth function of the $j$-th coordinate in $\mathbf{x}$, namely $x_{j}$. In various applications, however, nonparametric AMs based on main effects alone may not lead to a sufficiently rich representation for predicting the outcome of interest: including interaction terms can lead to better predictive models while remaining interpretable to a practitioner (Hastie and Tibshirani, 1987). AMs with pairwise interactions are useful in applied statistical modeling with various applications in medical sciences and healthcare, e-commerce applications, recommender system problems, and sentiment analysis, among others (Hastie, Tibshirani and Wainwright, 2015).
The need for interactions. Exploratory analysis on US Census Planning Database Tract dataset suggests that interaction effects across features can lead to improved predictions of the survey self-response rate. For example, there is interaction between the percentages of “People who do not speak English well” and “Renters” in the area. In areas with a relatively high concentration of poor English speakers (e.g., $\geq 5.4\%$), the self-response rate decreases on an average by $0.33\%$ for a unit increase in the percentage of renters. On the other hand, when the concentration of poor English speakers is relatively low (e.g., $\leq 1.6\%$), the self-response rate decreases at the rate of $0.21\%$. Similarly, there is a strong interaction effect between covariates “Single-unit households” and “Household moved in 2010 or later” in predicting the low self-response rate. Indeed, Erdman and Bates (2016) note the importance of incorporating interaction effects when predicting the self-response rate, although the paper does not pursue statistical modeling involving interactions.
@@ 2.2 Parsimonious models via $\ell_{0}$-penalization
Here, we study $\ell_{0}$-type estimators that limit the number of components in the additive models introduced earlier.
@@ 2.2.1 Sparse pairwise interactions
While a significant body of work has been devoted to studying sparsity in the context of linear models, sparsity in nonlinear models has received relatively less attention. Interestingly, we observe that the notion of parsimony is linked to the model being used and changes, for example, depending on whether we use a linear interaction model of the form $\mathbb{E}(y|\mathbf{x})=\sum_{j}x_{j}\beta_{j}+\sum_{j<k}x_{j}x_{k}\beta_{jk}$ or model (1) that has nonlinear components. For motivation, we consider Figure 3, which presents our findings on a 2019 US Census Bureau Planning Database dataset with $40$ covariates [^8] These covariates were motivated by our discussions with researchers at the US Census Bureau.. These features include covariates used for the low-response-score (Erdman and Bates, 2014), important covariates highlighted in Appendix C of the 2019 US Census Bureau Planning Database Documentation (US Census Bureau, 2019), plus some additional covariates capturing internet penetration and urbanization. Specifically, we observe that:
@@
  • The above nonparametric additive model leads to better test predictions than its linear model counterpart.
  • The best (based on validation tuning with prediction error) model with linear main and interaction effects contains $37$ main and $555$ interaction effects. On the other hand, using nonparametric AMs, we need 36 nonlinear main and interaction effects to obtain a similar test prediction performance. The nonlinear model also uses fewer covariates compared to the linear model counterpart.
The above observations suggest that the nonlinear model with sparse interactions has advantages over its linear model counterpart for predicting self-response rates. As discussed in Section 5, the above findings generally carry over to the expanded dataset with more features.
@@
Figure 3: Sparsity pattern of the main and interaction effects presented in a $p\times p$ matrix: a black square on the diagonal indicates the presence of a main effect, and an off-diagonal black square indicates the presence of an interaction effect in the joint model. [Left] Panel illustrates the sparsity pattern of a Lasso model with main and interaction effects. There are 37 main and 555 interaction effects in the optimal model. [Right] Panel illustrates the sparsity pattern of a nonlinear AM with main and interaction effects, i.e., model (4). There are 8 main and 92 interaction effects in the optimal model. Model (4) has prediction performance similar to the Lasso model, with only 3 main and 33 interaction effects. Nonlinear models lead to much more compact models and, hence, are easier to interpret than linear models with interactions. Both models were trained on a 2019 US Census Bureau Planning Database dataset (predicting the tract-level self-response rate) with $p=40$ covariates and $74,000$ observations.
@@
to simplify the expressions and propose an estimator based on the following problem:
@@
Related Work. In the special case of (4), when no interactions are present, several convex relaxations of the group $\ell_{0}$-penalty have been studied (Ravikumar et al., 2009; Zhao and Liu, 2012; Meier, van de Geer and Buhlmann, 2009; Buhlmann and van de Geer, 2011) [^11] In terms of existing implementations, R package SAM presents specialized algorithms for a convex relaxation of (4) without interactions. SAM, however, would not run on the dataset we consider here.. Hazimeh, Mazumder and Radchenko (2023) consider $\ell_{0}$-formulations for the setting without interaction effects and demonstrate the merits of using $\ell_{0}$-based formulations over convex relaxation-based approaches. Despite the appeal of estimator (4), computational challenges appear to be a key limiting factor in exploring this model in practice. In Section 2.4, we present a new algorithm for problem (4).
@@ 2.3 Statistical Theory
In this section, we explore the statistical error bounds of our proposed estimator in the deterministic design setting. The supplementary material (Ibrahim et al., 2024) provides technical details, assumptions, and proofs of the established results.
@@
Theorem 2.1 presented below derives a general non-asymptotic oracle prediction error bound, comparing the performance of our estimator to that of sparse approximations to the true regression function. Bounds of this type have been established for sparse nonparametric AMs with main effects; however, we are unaware of similar existing bounds for nonparametric models with pairwise interactions. The existing work has focused mainly on the Lasso-based estimators, which encourage sparsity in the main effects by penalizing the magnitudes of the functional components (see Meier, van de Geer and Buhlmann, 2009; Tan and Zhang, 2019, and the references therein). The use of $\ell_{1}$-based relaxations, in lieu of $\ell_{0}$-penalization, to induce sparsity in the main effects can lead to unwanted shrinkage, which may interfere with variable selection. The approach in Hazimeh, Mazumder and Radchenko (2023) directly controls the number of main effects, demonstrating the benefits of $\ell_{0}$-regularization theoretically and empirically.
Our analysis accounts for the interaction effects and focuses on the $\ell_{0}$-penalized formulation, which poses additional theoretical challenges relative to the $\ell_{0}$-constrained formulation. To our knowledge, the error bounds established here for sparse nonparametric AMs with pairwise interactions are novel.
@@
$\|\widehat{\bf f}_{n}-{\bf f}^{*}\|^{2}_{n}+\lambda_{n}\Omega_{\text{\rm gr}}(\widehat{f}_{n})+\mu_{n}G(\widehat{f}_{n})\lesssim\inf_{f\in\mathcal{C}_{\text{gr}}}\Big{[}\|{\bf f}-{\bf f}^{*}\|^{2}_{n}+\lambda_{n}\Omega_{\text{\rm gr}}(f)+\mu_{n}G(f)\Big{]}+\sigma^{2}\Big{[}r_{n}^{2}+\frac{\log(ep)}{n}\Big{]}$
@@
We note that a popular approach in the literature, even in the linear setting, has been to assume a sparse underlying model for the data. In contrast, Theorem 2.1 does not assume a ground truth sparse additive model, allowing for model misspecification and incorporating the approximation error into the bound. This feature is especially appealing for our application of interest where the ground truth is unavailable. The following corollary complements the general result in Theorem 2.1 by establishing a non-asymptotic prediction error bound for the proposed approach in the case where the additive model is correctly specified.
@@ 2.4 Efficient Computations at Scale
We present specialized algorithms for obtaining good solutions to Problem (4). Our approach scales to the problem-size of the Census application, with $n\approx 10^{5}$ and $p\approx 500$, which poses formidable computational challenges due to the presence of approximately $10^{5}$ interaction effects [^12] Note that, using 25 knots for every component, this leads to estimating around 2.5 million basis coefficients..
To obtain good solutions at scale, we use techniques inspired by first-order methods in continuous optimization (Nesterov, 2003) and careful exploitation of the problem structure. A high-level summary is presented below, with the details relegated to the Supplement (Ibrahim et al., 2024).
@@
State-of-the-art approaches. To appreciate the computational challenges of sparse nonlinear AMs with interactions, and nonparametric AMs in general, we provide a few examples of the problem instances that can be handled by current state-of-the-art algorithms with publicly available implementations. Current implementations based on R package SAM (Zhao and Liu, 2012), the stepwise GAM function in R package step.gam (which performs greedy variable selection), and Python package pyGAM, take on the order of days to run and/or face numerical difficulties for instances with $n\approx 10^{5}$ to obtain a single solution without interactions. pyGAM is not designed to do variable selection. Wood et al. (2017) present an interesting approach for AMs that scales to large $n$ settings (see R package mgcv) but doesn’t appear to perform automated variable selection in the presence of a large number of features – Wood et al. (2017) report instances containing fewer than $20$ pre-specified main and interaction effects. EBM (Nori et al., 2019) is a state-of-the-art computational approach for AMs, but does not allow for feature selection as all the main effects are included in the estimated model. Additionally, the variant of EBM that selects interaction effects results in many interaction effects, leading to sub-optimal support recovery – see Sections 4 and 5 for an illustration. The GAMI-Net method (Yang, Zhang and Sudjianto, 2021) is computationally expensive. Additionally, since it uses a multi-layered NN for every interaction effect it requires 10100×10-100$\times$10 - 100 × more parameters when compared with the approaches we consider.
@@
Algorithms for sparse nonlinear interactions: Problem (4). We represent the main and interaction effects as linear combinations of cubic spline basis functions (see the Supplement (Ibrahim et al., 2024) for the specific details). In particular, we let $\mathbf{f}_{j}=\mathbf{B}_{j}\boldsymbol{\beta}_{j}$, where $\mathbf{B}_{j}\in\mathbb{R}^{n\times K_{j}}$ is the model matrix and $\boldsymbol{\beta}_{j}\in\mathbb{R}^{K_{j}}$ is the vector of coefficients for each main effect component. Similarly, we let $\mathbf{f}_{j,k}=\mathbf{B}_{j,k}\boldsymbol{\theta}_{j,k}$, where $\mathbf{B}_{j,k}\in\mathbb{R}^{n\times K_{j,k}}$ is the model matrix and $\boldsymbol{\theta}_{j,k}\in\mathbb{R}^{K_{j,k}}$ is the vector of coefficients for each interaction effect component. Writing $\boldsymbol{\beta}$ for the vector obtained by stacking together the coefficients $\boldsymbol{\beta}_{j}$, $j\in[p]$ for the main-effects, and defining the vector $\boldsymbol{\theta}$ for the interaction effects analogously, we express the objective function in (2) as follows:
@@
Here, $\boldsymbol{S}_{j}=\boldsymbol{D}_{j}^{T}\boldsymbol{D}_{j}$ and $\boldsymbol{S}_{j,k}=(\boldsymbol{D}_{j}^{T}\boldsymbol{D}_{j})\otimes\boldsymbol{I}_{k}+\boldsymbol{I}_{j}\otimes(\boldsymbol{D}_{k}^{T}\boldsymbol{D}_{k})$ are the smoothness penalty matrices for the main effects and the interaction components, respectively (further details are provided in the Supplement (Ibrahim et al., 2024)). For convenience, we use the same smoothness penalty $\lambda_{1}$ for both the main and the interaction effects, though in general, they may be taken to be different. The above representation leads to the following form of the optimization problem (4):
@@
The indicator functions in the above equation are applied to vectors of basis coefficients corresponding to particular main or interaction effects. Hence, for example, when $\mathbb{1}\bigl{[}\boldsymbol{\beta}_{j}\neq\boldsymbol{0}\bigr{]}$ equals zero, the entire main effect $\mathbf{f}_{j}$ is also zero. As mentioned earlier, (8) can be expressed as a MIP and solved for small-to-moderate scale problems using modern MIP solvers. Developing specialized branch-and-bound based solvers for Problem (8) along the lines of Hazimeh, Mazumder and Saab (2022b); Hazimeh, Mazumder and Radchenko (2023) is beyond the scope of this work. We present fast approximate algorithms to obtain good solutions to Problem (8) for a family of tuning parameters. We note that the objective in Problem (8) is a sum of a smooth convex loss function and a discontinuous regularizer separable across the components $\{\boldsymbol{\beta}_{j}\}$ and $\{\boldsymbol{\theta}_{i,k}\}$. Motivated by the strong empirical performance of cyclical coordinate descent (CD) methods (Wright, 2015) in $\ell_{0}$-penalized linear regression (Hazimeh and Mazumder, 2020a), we explore block CD methods to obtain fast approximate solutions for the nonparametric setting with interactions (8). For convergence guarantees of this procedure, see Hazimeh and Mazumder (2020a) and references therein.
@@
The block minimization problems (9) and (10) can be solved in closed-form as discussed in the Supplement (Ibrahim et al., 2024). These block CD updates need to be paired with several computational devices in the form of active set updates, cached matrix factorizations, and warm-starts, among others. We draw inspiration from similar strategies used in CD-based procedures for sparse linear regression (Hazimeh and Mazumder, 2020a; Friedman, Hastie and Tibshirani, 2010), and extend them to our problem. See the Supplement (Ibrahim et al., 2024) for more details.
CD methods have old roots in optimization: see for example, the review paper Wright (2015). CD methods have close links with the well known Gauss-Seidel method (for solving linear/nonlinear systems of equations)—similar methods have been extensively used in additive models (Hastie and Tibshirani, 1990), where they are referred to as backfitting. CD schemes (including their block variants) arising from the optimization literature can be efficiently applied to optimization problems involving sparsity constraints: such problems arise in fitting additive models with sparsity-inducing penalties (Ravikumar et al., 2009; Hastie, Tibshirani and Wainwright, 2015; Hazimeh, Mazumder and Radchenko, 2023). We note that there are other successful additive model fitting approaches, for example, smooth backfitting (Mammen, Linton and Nielsen, 1999; Mammen and Park, 2006), which have excellent theoretical properties. These works focus on the additive model setup without sparsity—extending these approaches to our setting of large-scale sparse additive models in a computationally efficient fashion is an interesting direction for future research.